# RoCoFT: Efficient Finetuning of Large Language Models with Row-Column Updates

### v2

@@ Appendix E Additional Neural Tangent Kernel Results
Table 12: Relative difference in kernels (compared to full parameter $\mathbf{K}_{\theta}$) on single-sentence and sentence-pair tasks for 64-shot tasks
@@
(d) Full Parameter $\mathbf{K}_{\bm{\theta}}$ (SNLI)
(e) 1-Row $\mathbf{K}_{\bm{\theta}_{R}}$ (SNLI)
(f) 1-Column $\mathbf{K}_{\bm{\theta}_{C}}$ (SNLI)
(g) Full Parameter $\mathbf{K}_{\bm{\theta}}$ (RTE)
(h) 1-Row $\mathbf{K}_{\bm{\theta}_{R}}$ (RTE)
(i) 1-Column $\mathbf{K}_{\bm{\theta}_{C}}$ (RTE)
(j) Full Parameter $\mathbf{K}_{\bm{\theta}}$ (QQP)
(k) 1-Row $\mathbf{K}_{\bm{\theta}_{R}}$ (QQP)
(l) 1-Column $\mathbf{K}_{\bm{\theta}_{C}}$ (QQP)
Figure 7: Neural Tangent Kernels on 16-shot training data for different tasks
Appendix F Dataset Description
@@
Appendix G Evaluation Metrics
@@
$\text{MCC}=\frac{\text{TP}\times\text{TN}-\text{FP}\times\text{FN}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}}.$ (2)

### v3

@@ Appendix G Additional Neural Tangent Kernel Results
Table 13: Relative difference in kernels (compared to full parameter $\mathbf{K}_{\theta}$) on single-sentence and sentence-pair tasks for 64-shot taskstab:kernel_diff_single_k64
@@
(d) $\mathbf{K}_{{}_{\text{LoRA}}}$ with $r=1$
(e) Full Parameter $\mathbf{K}_{\bm{\theta}}$ (RTE)
(f) 1-Row $\mathbf{K}_{\bm{\theta}_{R}}$ (RTE)
(g) 1-Column $\mathbf{K}_{\bm{\theta}_{C}}$ (RTE)
(h) $\mathbf{K}_{{}_{\text{LoRA}}}$ with $r=1$
(i) Full Parameter $\mathbf{K}_{\bm{\theta}}$ (QQP)
(j) 1-Row $\mathbf{K}_{\bm{\theta}_{R}}$ (QQP)
(k) 1-Column $\mathbf{K}_{\bm{\theta}_{C}}$ (QQP)
(l) $\mathbf{K}_{{}_{\text{LoRA}}}$ with $r=1$
(m) Neural Tangent Kernels on 16-shot training data for different tasks
Appendix H Dataset Description
@@
Appendix I Evaluation Metrics
@@
[…]
$\frac{\text{TP}\times\text{TN}-\text{FP}\times\text{FN}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}}.$ (5)